LEP EXPERIMENTS COMMITTEE

Minutes of the 37th meeting,
2 June 1994






OPEN SESSION
 

1. Start-up and performance of LEP 1

S. Myers reported on the rapid LEP start-up. LEP has reached a peak luminosity of 1.5 x 1031 cm-2s-1 and an integrated luminosity of 500 nb-1 per day. The target for 1994 is 60 pb-1 to be compared with last yearís total of 40 pb-1).

2. Progress report on the LEP 2 superconducting rf system

D. Boussard presented a progress report on the work of the SL-RF group in preparing the LEP 2 sc rf cavities. The areas of particular concern are cavity production, power couplers, higher-order-mode couplers, and in-situ recovery of cavity performance. A sc rf module 4 cavities) will be installed in the floating 2-week technical stop, which is foreseen for September. During the October technical stop, a second sc rf module will be installed, with improved couplers of perhaps the final design. If machine tests on the latter module prove successful, production retro-fitting could then proceed on the full set of sc rf cavities for LEP 2.

3. Status report on the bunch-trains working group

T. Camporesi presented a status report on the bunch-trains working group. After considering the constraints of the machine and the detectors, the plan is to develop LEP operation with 4 x n bunches per beam where the number of bunches per train, n = 2, 3 or 4) and with a maximum bunch-train width of about 750 ns. This scheme would involve head-on crossings and vertical beam separation near the interaction regions to avoid parasitic beam collisions. Recent tests have confirmed that large bumps do not create aperture or injection problems for the machine nor do they produce serious backgrounds in the detectors. If the bunch-trains scheme proves successful, LEP could potentially produce 120 pb-1 in 1995 by operating with 4 x 4 bunches per beam to be compared with a total since LEP started of about 90 pb-1 per detector, up to the end of 1993). A final document from the working group is expected in August.

4. Reports on the LEP experiments

OPAL:     R.D. Heuer
DELPHI: J.E. Augustin
ALEPH:   L. Foà
L3:           V. Innocenti
Several of the LEP detectors especially ALEPH and DELPHI) reported present beam backgrounds substantially larger than in 1993, but significantly improving as the cycle continues. OPAL presented an upgrade proposal involving a layer of plastic scintillator located directly in front of the end-cap electromagnetic calorimeter and read out via wave shifter fibres into multi-anode phototubes. ALEPH reported that their superconducting solenoid had been successfully repaired in an impressive operation. L3 requested that the 94/95 shutdown be extended until the end of May 1995 in order to complete the detector upgrade and prepare for multi-bunch operation.
 

CLOSED SESSION

Present: J. Allaby, W. Bartel, L.L. Camilleri, T. Camporesi, P. Darriulat, J.D. Dowell Chairman), R. Eichler, J. Feltesse, L. Foà,
B. Foster, W. Hollik, K. Hübner, P.G. Innocenti, J. Kirkby Secretary), K.H. Kissler, G. Matthiae, A. Menzione, D.J. Miller, S. Myers,
L. Rolandi, D. Schlatter, A. Weidberg, D.O. Williams and C. Wyss.

Part time: A. Blondel, D. Boussard and C.H. Llewellyn Smith.

Apologies: G. Coignet, G. Kantardjian, H. Wenninger, P.M. Zerwas and F. Zwirner.
 

1. Approval of the minutes of the 36th meeting

The minutes of the 36th meeting LEPC 94-6, LEPC 36) were approved without modifications.
 

2. Chairman's report

The Chairman joined with the committee in congratulating the LEP team for the rapid machine start-up this year.

It was reported that the Research Board had approved all the recommendations made at the last LEPC meeting.
 

3. Discussion on LEP 2 progress

C. Wyss presented an update on aspects of the LEP 2 programme not covered by the talk of D. Boussard in the open session. Preparations for all of the systems-cryogenics, RF power supplies, controls, vacuum chambers, superconducting quadrupoles, vertical separators and collimators-are proceeding smoothly. The working plan is to install at least 8 sc rf modules during the 94/95 shutdown, with the aim of running at 45 GeV in 1995 with sc cavities only.
 

4. Discussion of bunch-train operation in LEP

Assuming the bunch train scheme is successful, the present peak luminosity of about 1.6 x 1031 cm-2s-1 could be raised by up to a factor of three, to 5 x 1031 cm-2s-1 with 4 x 4 bunches per beam). In the case of LEP 2, the peak luminosity could reach 7 x 1031 cm-2s-1 with 4 x 2 bunches per beam), which is five times the design luminosity. The committee noted that all four LEP detectors have indicated they can operate up to a maximum train width of about 750 ns corresponding to 4 x 4 bunches) - and that the upgrade costs are reasonable 200-300 kCHF per detector). However, ALEPH has indicated that their data suffer some deterioration beyond 600 ns train width and, although accepting 4 x 4 bunches at LEP 1, they would oppose trains longer than 600 ns at LEP 2. Also noted were the positive results from the bunch-train tests that have been made so far, although definitive tests cannot take place until after the October technical stop. The committee therefore expressed strong support for LEP to operate with bunch trains, commencing at the beginning of 1995, and encouraged the LEP detectors to make the necessary preparations.
 

5. Discussion of the L3 request to extend the 94/95 shutdown

The LEPC indicated a strong reluctance towards an extension of the 94/95 shutdown until the end of May, which was requested by L3 in the open session. The next year is seen both as an important year for physics high-luminosity LEP 1 operation with bunch trains) and as a crucial year for LEP 2 preparations. L3 is therefore encouraged to adjust its plans for the 94/95 shutdown to fit within the presently-foreseen schedule. If the L3 collaboration wishes to make a formal request for this extension, a detailed paper would be required for the next meeting of the LEPC.

6. Discussion of the Proposal LEPC 94-8/P8: A Test of Longitudinal Spin Polarization at LEP

Since there was no open presentation of this proposal, A. Blondel was invited to present a summary in the closed session. This was preceded by a brief presentation from W. Bartel of the impressive new results obtained with the spin rotator at HERA. A transverse polarization of 65% has been measured for the electron beam at 27.5 GeV, and it is not affected by turning on the spin rotator or by beam collisions. Successful longitudinal spin polarization has therefore been achieved. After a discussion of the LEP proposal and the severe constraints on the LEP programme, the proponents were invited to make an open presentation at the next meeting, and three referees W. Bartel, W. Hollik and G. Matthiae) were appointed.
 

7. Discussion of the Proposal LEPC 94-9/P3: Proposal to Install Scintillating Tiles in the OPAL Endcap

In view of the late receipt of this paper, the committee was unable to reach a conclusion at the present meeting. The proponents are invited to present the upgrade proposal at the next open session. In addition, a referee B. Foster) was appointed and asked to make a report at the same meeting.
 

8. Discussion of the Memorandum LEPC 94-10/M109: Search for Cosmic Ray Coincidences with Detectors in the LEP tunnel

The committee took note of this memorandum, which is ultimately aimed at a search for cosmic ray coincidences in several detectors around the LEP ring. The present proposal involves only the ALEPH detector and scintillator arrays placed nearby in the LEP tunnel. The committee requested a presentation be made at the next open session.
 

9. LEP schedule for 1994

The likely dates for the ëfloatingí 2-week technical stop for installation of a superconducting rf module are in September 1994. The precise dates will be fixed shortly by the LEP co-ordinator, in consultation with the LEP accelerator team, and communicated to the LEP detectors.
 

10. Any other business

The committee requested that G. Von Holtey present a progress report at the next open session on the LEP 2 detector masks. Of particular interest are the results from the prototype LEP 2 masks in DELPHI, their compatibility with bunch-trains operation, and the plans for equipping all detectors.

The Chairman reported that the dates for the Cogne meeting to review LEP physics and plans are 25-30 September 1995.

Finally, the Chairman thanked the outgoing LEPC members -L.L. Camilleri, R. Eichler, J. Feltesse, A. Menzione and P. Zerwas- for their excellent work on the committee, and warmly expressed his gratitude to the outgoing Research Director, Pierre Darriulat, for his continuous advice and support.
 

11. LEPC meetings in 1994

The date of the next meeting of the LEPC will be Friday 16 September. In order to better match the expected progress of LEP, the final LEPC meeting of 1994 has been postponed from the original dates of 7-8 November to Friday 9 December.
 
 

                                                                                                                             J. Kirkby
 

DOCUMENTS RECEIVED

[1] Proposal: A Test of Longitudinal Spin Polarization at LEP G. Alexander et al.), LEPC 94-8/P8.

[2] Proposal to Install Scintillating Tiles in the OPAL Endcap OPAL Collaboration), LEPC 94-9/P3.

[3] Search for Cosmic Ray Coincidences with Detectors in the LEP tunnel A. Ball et al.), LEPC 94-10/M109.